The sum of the measures of a complement and a supplement of an angle is 200. What is the measure of the angle?
Answer by rfer(10417)

(180-x)+(90-x)=200
180-x+90-x=200
-2x=70
x=-35
145+55=200
145 degrees
55 degrees